OpenStudy (anonymous):

Solve, How long will it take for the population of a certain country to double if its annual growth rate is 3.1% ? round to the nearest year

OpenStudy (kropot72):

Let the current population be N and let the number of years for the population to double be x. Then we can express the growth in the following equation: \[2N=N(1+0.031)^{x}\ .......(1)\] Equation (1) can be rearranged as follows: \[(1.031)^{x}=2\ .........(2)\] Taking natural logs of both sides of equation (2) we get \[x\times \ln 1.031=\ln 2\ ........(3)\] Dividing both sides of equation (3) by ln 1.031 we get an expression for x as follows: \[x=\frac{\ln 2}{\ln 1.031}=you\ can\ calculate\] Hint: Remember to round your result to the nearest year.

OpenStudy (anonymous):

How can i round it to the nearest year

OpenStudy (anonymous):

well all you have to do is plug it into your calculator and do the numbers that are to the left of the decimal place